### v1

# FREE-Merging: Fourier Transform for Model Merging with Lightweight Experts

Abstract
In the current era of rapid expansion in model scale, there is an increasing availability of open-source model weights for various tasks. However, the capabilities of a single fine-tuned model often fall short of meeting diverse deployment needs. Model merging has thus emerged as a widely focused method for efficiently building a single model tailored for multiple tasks combined from existing models. Nevertheless, existing model merging methods face challenging trade-offs between performance and deployment costs, primarily due to task conflicts within the merged network. Our analysis of neural networks reveals that some task-specific information introduced by fine-tuning minimally enhances performance but heavily impacts generalization, leading to task conflicts. To mitigate the impact of this information, we propose FR-Merging, an innovative method that leverages frequency domain information to efficiently filter harmful specialized information, thereby minimizing the impact of task conflicts on the backbone with minimal cost. Since performance loss is inevitable with cost-free merging methods, we introduce a lightweight task-specific expert that can be dynamically integrated during inference to compensate for information loss. This framework, FREE-Merging (FR-Merging with lightweight experts), strikes a balanced trade-off between training cost, inference speed, storage requirements, and performance. We demonstrate the effectiveness of both FR-Merging and FREE-Merging on multiple tasks across CV, NLP, and Multi-Modal domains and show that they can be flexibly adapted to meet specific needs.
1 Introduction
With the rapid development of deep learning, the pretrain-finetune paradigm has become the standard methodology [32, 8, 35]. However, in many edge-side applications, users often lack powerful computing resources and large datasets, preferring ready-to-use models tailored to specific scenarios. The recent growth of open-source platforms like HuggingFace [38] has made this more feasible, allowing easy access to model weights for various tasks. Yet, practical demands are often not fully covered by a single existing model and are typically a union of several models [46, 19]. However, if each subtask is handled by different models, it will raise large storage and inference costs in the era of large models. Model merging attempts to directly combine existing model weights to build a model capable of addressing all target tasks [18, 28, 42]. This method mitigates the high training costs and data privacy issues faced by multi-task learning (MTL) [49, 17], reducing both training and deployment costs, and has garnered widespread attention.
However, as model merging seeks to minimize training costs and data, it inevitably encounters performance degradation due to task conflicts [17]. Current solutions fall into three categories. The first method calculates the merging coefficients based on weights or data [19, 28]. The second method uses the parameter differences before and after fine-tuning (task vectors) as the merging target [18, 42, 46]. The third method introduces independent knowledge for each task [17, 24]. Although the third way improves the balance between performance and deployment costs, it faces two issues. On the one hand, it overlooks backbone network optimization during merging, leading to subpar performance. On the other hand, it demands great computation and storage, making it unsuitable for edge deployment.
We believe that developing an effective model merging method requires addressing both of the above issues. On the one hand, since the backbone comprises the majority of parameters and thus determines task performance, an efficient merging mechanism is needed to reduce performance loss. On the other hand, based on the no free lunch theorem, cost-free model merging cannot avoid task conflicts. Introducing task-specific experts is unavoidable. However, it is crucial to ensure that these experts do not significantly burden storage or inference. Therefore, we propose our two-stage optimization method, FREE-Merging, as shown in Fig 1. It provides an efficient backbone and lightweight experts.
Figure 1: The workflow of FREE-Merging involves two main steps. First, FR-Merging utilizes high-pass filtering to remove harmful specialized information from each model, thereby constructing a high-quality merged backbone. Second, it employs lightweight extraction of task experts, which are dynamically added during inference to mitigate the impact of task conflicts.
For the optimization of backbone merging, we propose that fine-tuning specializes parameters for specific tasks [9], so the fine-tuned parameters carry task-specific information. Task-specific information leads to strong task characteristics, which reduces generalization ability and causes task conflicts during merging [45]. Therefore, we aim to mitigate the impact of this information on generalization without affecting model performance, thereby reducing performance loss during merging. Identifying task-specific information in the original parameters has long been challenging [29]. Naturally, the frequency domain comes into our eyes. After the Fourier transform, high-frequency signals represent sharp changes in the signal, while low-frequency components capture the overall structure. Thus, the low-frequency part is more likely to contain the unique aspects of the signal, in this case, the information most relevant to the task. We demonstrate in Sec 4.2 that high-pass filtering of parameters significantly enhances the generalization ability with minimal impact on performance, thus avoiding performance loss due to task conflicts during merging. This computationally lightweight yet highly effective method, FR-Merging, is the first to demonstrate the efficacy of Fourier filtering on neural network parameters, showing potential for broader deep learning applications.
After obtaining a high-performance merged backbone, we utilize a lightweight task-specific expert to compensate for the inevitable information loss during merging. Our goal is to retain task-relevant information with minimal storage. Unlike existing methods that require extensive knowledge storage, our method focuses on rescaling low-frequency signals, which store task-specific information. This allows us to maintain task expertise with only 1% of the parameters. Additionally, to improve inference speed, we further simplify the expert. Inspired by MoE [34], we use a router for dynamic routing of experts, enhancing model flexibility.
In summary, this paper proposes FREE-Merging, which utilizes FR-Merging to enhance model generalization and uses lightweight experts to mitigate information loss of model merging. The main contributions are: 1). We identify the relationship between the low-frequency region of fine-tuning parameters and model generalization, which helps to reduce task conflicts during merging and has significant potential in deep learning. 2). We theoretically validate the necessity of introducing new information in merging and propose an effective lightweight expert construction method to compensate for information loss during merging. 3). According to extensive experiments, the proposed FR-Merging and FREE-Merging demonstrate their effectiveness across various tasks in vision, language, and multi-modal domains.
2 Related Work
Model Merging. Model merging combines existing model weights to handle multiple tasks without training [18, 19, 43]. The main challenge lies in resolving task conflicts [42]. Simple averaging [39] often leads to significant performance degradation. Methods like Fisher-Merging [28] and RegMean [19] employ different matrices to determine merging coefficients. However, they demand considerable computational costs or access to data, making them unsuitable for edge deployment where both computational power and data are limited. Task Arithmetic [18] merges task vectors instead of weights, while Ties-Merging [42] addresses parameter conflicts, and AdaMerging [46] automates the setup of merging coefficients. Nonetheless, these methods require training and have limited applicability. Techniques such as DARE [47] and PCB-Merging [11] attempt to alleviate performance drops by analyzing parameters, yet they still encounter task conflicts. EMR-Merging [17] and Twin-Merging [24] propose storing knowledge for each task, but they require substantial storage and do not address task conflicts within the merged backbone. Our method uses high-passing filtering to reduce task conflicts in the backbone while utilizing lightweight experts to enhance performance. We achieve a balance between storage, inference speed, and performance, showing impressive results across various tasks. For further details, please refer to Appendix 9.2.
Model Ensemble. Model ensemble combines outputs from multiple models [14, 10], but with large models, it faces storage and inference challenges. In contrast, Model merging enables a single model to perform multiple tasks, greatly reducing storage and inference costs.
Multi-task Learning. Multi-task learning (MTL) aims to solve multiple tasks using a single model [4, 36, 23], but in the era of large models, it faces challenges like high training costs, data privacy issues, and expertise requirements [46]. In contrast, Model merging uses existing open-source model weights to create a multi-task model with little or no training, significantly reducing deployment costs.
3 Preliminaries
Notation. Assume that $f_{\theta}(x)\rightarrow y$ represents a neural network parameterized by $\theta$, where the input $x\in R^{m}$ and the output $y\in R^{n}$. Recent studies have focused on model merging via task vectors. For task $k$, the task vector is defined as $v_{k}=\theta_{k}-\theta_{pre}$, where $\theta_{k}$ is the fine-tuned model for task $k$, and $\theta_{pre}$ is the corresponding pre-trained model. Additionally, $e_{k}$ represents the task expert that retains the ability to solve task $k$ with few parameters.
Problem Definition. The task of model merging aims to combine models $\{\theta_{k}\}_{k=1}^{K}$ for tasks $T=\{t_{i}\}_{i=1}^{K}$ into a single model $\theta_{m}$ that can handle all tasks in $T$. This paper focuses on a more practical scenario, where a shared pre-trained model $\theta_{pre}$ serves as the base for all fine-tuned models. It is formulated as $\theta_{m}=\lambda\sum_{k=1}^{K}\hat{G}(\theta_{k})$, and when based on the task vector, it becomes $\theta_{m}=\theta_{pre}+\lambda\Sigma_{k=1}^{K}G(v_{k})$, where $\lambda$ is the coefficient, and $\hat{G}/G$ is transformation on the parameters. Assuming a consistent pre-trained model aligns with trends toward foundation architectures, like LLaMa [35] for language and CLIP [32] for vision, which are then fine-tuned for specific tasks.
4 Methodology
This section presents the proposed FREE-Merging: Sec 4.1 explains the need for a two-part design, Sec 4.2 and Sec 4.3 cover the FR-Merging for the backbone and the lightweight expert module, respectively, followed by the overall FREE-Merging framework and a complexity analysis in Sec 4.5.
4.1 Motivation
In this section, we discuss why it is necessary to propose an optimized merging method for the backbone and introduce the expert module during inference, respectively.
First, merging optimization for the backbone is crucial, yet often overlooked by current expert-based model merging methods. As the backbone holds most parameters, its performance is key to the overall success. For instance, upgrading from LLaMa2 [35] to LLaMa3 [12] significantly improves results even with the same fine-tuning methods and data. Thus, we propose a training-free backbone merging method for reducing task conflicts in the backbone.
Next, we present a theorem showing the necessity of task experts. Based on Theorem 4.1, introducing additional information is essential. For proof, see Appendix 8, and the effectiveness of task experts is proved in Appendix 10.
Theorem 4.1
Model Merging has no free lunch. Merged model $\theta_{m}$ cannot simultaneously retain the capabilities of $\{\theta_{k}\}_{k=1}^{K}$ without introducing additional information.
4.2 FR-Merging
In this section, we design an efficient approach to reduce task conflicts in response to the backbone optimization discussed in Sec 4.1. This requires further analysis of the fine-tuning. Fine-tuning can be viewed as mapping the general representation space of a pre-trained model to the space of a specific task [44]. Since different fine-tuning weights fall at different positions in the loss landscape of a task, linearly interpolating the weights could place them in a high-loss region, which is known as task conflict [1]. Therefore, we aim to increase the overlap of the positions of each model weight within the loss landscape of the same task, so that merged models are more likely to fall into a loss basin. Variations in the loss landscape arise from task-specific characteristics in the parameters [40], so we need to eliminate task-specificity in the parameters while preserving the task performance.
If we consider neural network parameters as signals carrying training information, inspired by image processing [31, 3], we transform the task vector into the frequency domain, where high-frequency signals represent sharp changes, and low-frequency components define the overall structure [5]. Since most of the information in the signal is contained within the overall structure, filtering out the low-frequency signals can be seen as removing the most distinctive set of task information. As a result, the likelihood that the linear interpolation falls within the loss basin increases [1]. We use $v(x,y)$ to represent the task vector and $G(x,y)$ to denote the transformed result. Then we have:
$G(x,y)=\mathcal{F}^{-1}\{H(\eta,\gamma)\cdot\mathcal{F}\{v(x,y)\}\},$ (1)
where
$H(\eta,\gamma)$ =$=\begin{cases}1\quad\sqrt{\eta^{2}+\gamma^{2}}\geq D_{0}\\ 0\quad\sqrt{\eta^{2}+\gamma^{2}}<D_{0}\end{cases},$ (2)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ represent the Fourier transform and inverse Fourier transform, respectively. $D_{0}$ denotes the cutoff frequency, which can be adjusted as a hyper-parameter.
Figure 2: High-pass filtering improves performance over the unfiltered version. Rows represent the added task vectors, and columns indicate the test datasets. The results show a significant boost in generalization with only a slight decline in task performance.
Experimentally, we validate that high-pass filtering significantly improves model generalization with minimal impact on model performance as shown in Fig 2. We fine-tune ViT-B/16 [32] on 30 visual tasks following [17]. After removing low-frequency information, the performance drop on the corresponding tasks (diagonal entries) is substantially smaller than the improvement in generalization (off-diagonal entries). This is a notable finding that similar to image processing, the low-frequency signals in the task vector carry more task-specific information that enhances performance on particular tasks while substantially reducing generalization. Therefore, applying filtering as FR-Merging during model merging can help alleviate task conflicts.
To ensure that the mean of the outputs remains consistent after model merging, we define merging coefficient $\lambda_{i}$ corresponding to the task vector $v_{i}$ for the task $t_{i}$ as:
$\lambda_{i}=\mathbb{E}(v_{i})(\sum_{j=1}^{K}\mathbb{E}(v_{j}))^{-1},$ (3)
where $\mathbb{E}$ denotes the average. Overall, we construct a merged backbone in this part that minimizes task conflicts.
4.3 Lightweight Expert Extraction
Since low-frequency signals are discarded in the backbone and information is lost during merging, this section constructs lightweight task experts to be added during inference, compensating for the task-specific capabilities. As analyzed in Sec  4.2, low-frequency signals contain task-specific information, but preserving them requires an inverse Fourier transform during each inference, which is inefficient. Thus, we propose a more efficient approach.
Since fine-tuning leads to task specialization, the most changed parameters during fine-tuning can be approximated as having the greatest impact on this process [27]. After reasonable rescaling, these selected parameters serve as task experts, minimizing storage (typically around 1% of parameters) and adding no extra computation during inference. The approximation’s validity is detailed in the Appendix 10.
Although previous work has proposed magnitude-based extraction [42], insufficient research on rescaling limits compression. Rescaling is crucial for expert extraction, so we propose a rescaling method. Let $M(v_{i},d)$ represent the top-$d$ parameters for task vector $v_{i}$, $\mu_{i}$ the rescale factor, and $e(v_{i})$ represents the expert for task $t_{i}$. Then we have:
$e(v_{i})=\mu_{i}M(v_{i},d),\ \mu_{i}=-\frac{\mathbb{E}(M(v_{i},d))\cdot\log(d)}{\lambda_{i}\cdot\mathbb{E}(v_{i})}.$ (4)
The goal is to maintain a consistent output after extraction while ensuring tasks with smaller backbone weights receive greater expert compensation to preserve performance.
In summary, we use an efficient method to extract task-specific experts, which can be integrated into the backbone to compensate for performance decline due to task conflicts without bring excessive storage and inference overhead.
4.4 FREE-Merging
Algorithm 1 Workflow of FREE-Merging
0:  Task vectors $\{v_{i}\}_{i=1}^{K}$, pre-trained model $\theta_{pre}$, router $R$, and input data $X$.
1:  FR-Merging: $\triangleright$ Only excute once.
2:  Get backbone $\theta_{m}=\theta_{pre}+\sum_{i=1}^{K}\lambda_{i}G(v_{i})$, where $G$ and $\lambda$ are defined by Eq. 1 and Eq. 3, respectively.
3:  Expert Extraction: $\triangleright$ Only excute once.
4:  Get task experts $\{e_{i}\}_{i=1}^{K}$ as Eq. 4.
5:  Inference:
6:  for $x\in X$ do
7:     $[w_{1},...w_{K}]\leftarrow argmax(R(x))$
8:     $\theta_{*}=\theta_{m}+\sum_{i=1}^{K}w_{i}e_{i}$
9:     $Y\leftarrow Y\cup f(x;\theta_{*})$
10:  end for
10:  Output $Y$ for input $X$.
In this section, we describe how to combine the two parts into FREE-Merging, making it suitable for multiple tasks without affecting interference speed. Inspired by MoE, we use a router to dynamically assign tasks to different activated parameters. The router can take various forms [20], allowing for more flexibility in handling different tasks. For detailed algorithms, see Alg 1.
First, we obtain the merged backbone (line 2) and the task experts (line 4). Then, during the inference, we dynamically route the experts based on inputs (lines 7-9).
Method Base Model SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
B L B L B L B L B L B L B L B L B L
Individual 75.3 82.3 77.7 92.4 96.1 97.4 99.7 100 97.5 98.1 98.7 99.2 99.7 99.7 79.4 84.1 90.5 94.2
Traditional MTL 73.9 80.8 74.4 90.6 93.9 96.3 98.2 96.3 95.8 97.6 98.9 99.1 99.5 99.6 77.9 84.4 88.9 93.5
Weight Averaging 65.3 72.1 63.4 81.6 71.4 82.6 71.7 91.9 64.2 78.2 52.8 70.7 87.5 97.1 50.1 62.8 65.8 79.6
Fisher Merging [28] 68.6 69.2 69.2 88.6 70.7 87.5 66.4 93.5 72.9 80.6 51.1 74.8 87.9 93.3 59.9 70.0 68.3 82.2
Task Arithmetic [18] 63.8 74.1 62.1 82.1 72.0 86.7 77.6 93.8 74.4 87.9 65.1 86.8 94.0 98.9 52.2 65.6 70.1 84.5
Ties-Merging [42] 64.8 76.5 62.9 85.0 74.3 89.3 78.9 95.7 83.1 90.3 71.4 83.3 97.6 99.0 56.2 68.8 73.6 86.0
Breadcrumbs [6] 63.0 74.9 61.0 84.9 75.7 88.6 84.7 95.4 83.4 89.8 76.4 90.1 98.1 99.2 57.9 68.2 75.0 86.4
PCB-Merging [11] 65.5 75.8 64.1 86.0 78.1 88.6 80.2 96.0 84.7 88.0 77.1 90.9 98.0 99.1 58.4 70.0 75.8 86.9
FR-Merging(ours) 66.2 76.4 64.5 87.0 77.2 90.2 90.1 96.8 85.4 92.0 82.3 92.8 98.5 99.3 60.0 71.5 78.1 88.3
AdaMerging++ [46] 66.6 79.4 68.3 90.3 82.2 91.6 94.2 97.4 89.6 93.4 89.0 97.5 98.3 99.0 60.6 79.2 81.1 91.0
Twin-Merging [24] 73.6 82.6 71.7 90.3 92.1 94.9 99.3 99.4 95.3 96.7 97.2 98.0 99.1 99.4 74.0 80.1 87.8 92.7
EMR-Merging [17] 74.1 82.3 72.7 90.8 91.9 95.3 99.4 99.5 95.8 96.9 96.9 98.1 99.1 99.5 72.1 80.1 87.7 92.8
FREE-Merging(ours) 77.1 83.5 78.2 92.4 93.4 96.2 99.5 99.6 96.3 98.1 98.2 98.7 99.5 99.7 75.4 81.7 89.7 93.7
Table 1: Comparison of model merging performance using ViT-B/32 (B) and ViT-L/14 (L) on eight visual tasks evaluating by accuracy.
4.5 Time Complexity and Storage Space Analysis
In this section, we analyze the time complexity of merging and the storage space required after merging. We aim to keep both as small as possible to meet practical user needs. Assume that there are $n$ models to merge, each with $m$ parameters and requiring storage space $s$.
The time complexity analysis can be divided into two parts: merging and inference. In the merging process, the time complexity of FR-Merging is $O(nmlogm)$ [13] while that of extracting the expert is $O(nm)$. Thus, the overall time complexity of merging is $O(nmlogm)$, which is manageable. For example, merging three 10B models requires only $O(10^{12})$ addition operations, which is not a large load for current devices. Additionally, since only a lightweight router is added, the extra inference cost is negligible.
Next, we analyze storage space. We typically save only about 1% additional parameters per model. Therefore, our storage space requirement is $(1+0.01n)s$, which is a significant improvement compared to the requirement of $ns$ for the model ensemble. For large models, even a 1% reduction in data storage can greatly alleviate storage pressure. For detailed analysis, please refer to Sec 5.7.
5 Experiments
In this section, we evaluate the proposed method across various tasks. We compare FR-Merging with other cost-free methods (which do not require additional training or data), including Weight Averaging, Fisher Merging [28], Task Arithmetic [18], Ties-Merging [42], Breadcrumbs [6], and PCB-Merging [11]. We also compare FREE-Merging with other popular methods, including RegMean [19], AdaMerging [46], EMR-Merging [17], and Twin-Merging [24]. The original EMR-merging uses a perfect router; however, we use an imperfect router for realism. Some experiments include Traditional MTL with training data from all tasks as a baseline. The evaluation covers CV, NLP, and multi-modal tasks, followed by ablation studies and further analysis.
5.1 Merging Vision Models
Merging 8 ViTs. We select the two visual encoders of CLIP, ViT-B/32 and ViT-L/14 [32], as the pre-trained models and choose eight image classification tasks as target tasks: SUN397, Cars, RESISC45, EuroSAT, SVHN, GTSRB, MNIST, and DTD [46]. Additionally, an MLP is used as the lightweight router. Detailed information on the datasets and baselines can be found in the Appendix 9.1.
The comparison results of various methods can be found in Table 1. First, compared to existing methods that do not require additional training and data storage, our proposed FR-Merging shows significant improvements, achieving approximately 2.5% on ViT-B/32 and around 1.5% on ViT-L/14 over the current state-of-the-art, PCB-Merging [11]. This indicates that the Fourier filtering effectively removes task-specific parameters from the task vector, greatly enhancing the generalization ability of the model and reducing performance loss due to task conflicts. Through FR-Merging, we can effectively construct a merged backbone that can be used independently at a low cost. Second, compared to methods requiring extra information, FREE-Merging also shows a considerable improvement about 2%. This advantage is based on the lightweight computation and efficient expert extraction of FREE-Merging, as we achieved this with 1% additional storage for each model, outperforming 3% of EMR-Merging [17] and 2% of Twin-Merging [24]. Thus, FREE-Merging provides an excellent balance between storage, computation, and performance.
Method Additional Cost Avg. Acc
Individual 93.02
Weight Averaging 42.51
Task Arithmetic [18] 48.88
Ties-Merging [42] 37.53
Breadcrumbs [6] 43.57
FR-Merging(ours) 53.90
RegMean [19] 68.13
AdaMerging [46] 60.24
Twin-Merging [24] 75.52
EMR-Merging [17] 76.39
FREE-Merging(ours) 79.67
Table 2: Comparison of model merging results using ViT-B/16 on 30 visual tasks. Additional Cost indicates whether additional training or additional data is required during model merging.
Merging 30 ViTs. We select ViT-B/16 as the pre-trained model and evaluate the model merging performance across 30 image classification tasks. The dataset selection and configuration of fine-tuning tasks follow [17].
The results are shown in Table 2. Details can be found in Appendix 11.1. It can be found that, compared to current methods, our proposed FR-Merging and FREE-Merging show remarkable performance. Although real applications rarely involve merging such a large number of models, we introduce this experiment mainly to demonstrate that the proposed Fourier filtering and lightweight task experts can mitigate performance loss caused by task conflicts even in extreme scenarios. This further illustrates the effectiveness and broad applicability of our methods.
5.2 Merging Language Models
Merging Medium-sized Language Models. We investigate the generalization ability of our method on medium-scale language models, using RoBERTa [50] as the pre-trained model with full fine-tuning on eight discriminative tasks [42]. Details can be found in Appendix 9.1.
The average results of the eight tasks after merging are shown in Table 3, and the complete results are available in Appendix 11. It can be observed that among our proposed methods, FR-Merging shows a significant advantage of nearly 2% over other cost-free methods. Additionally, FREE-Merging also exhibits remarkable effectiveness, achieving optimal results on almost all datasets and outperforming current methods in average performance. This demonstrates the applicability and great potential of our proposed methods for language models.
Merging PEFT models. PEFT is a popular method that reduces fine-tuning costs [9], but storing PEFT parameters for multiple tasks remains space-inefficient [48]. The effectiveness of our approach on PEFT demonstrates the generalization capability and offers an efficient deployment way.
Firstly, we use T0-3B [33] as the pre-trained model and $(IA)^{3}$ [22] as the PEFT method, fine-tuning on 11 discriminative tasks [17]. Our proposed FR-Merging and FREE-Merging show great advantages with an improvement of nearly 1% in Table 3. Then, we select Qwen-14B [2] as the pre-trained model and LoRA [16] as the PEFT way, fine-tuning on three generative tasks before merging [24]. The results show that our method exhibits nearly a 2% advantage, whether in cost-free conditions or the complete framework. This demonstrates that our methods remain highly feasible even with PEFT, confirming their generalization ability. Full results are available in the Appendix 11.
Method Base Model Full PEFT Avg.
RoBERTa T0-3B Qwen-14B
Individual 85.55 71.35 72.07 76.32
Weight Averaging 51.34 58.01 66.80 58.71
Task Arithmetic[18] 66.65 63.91 66.40 65.65
Ties-Merging [42] 64.04 66.41 67.46 65.97
Breadcrumbs [6] 68.19 54.64 66.77 63.20
PCB-Merging [11] 67.86 66.10 66.69 66.39
FR-Merging(ours) 70.02 66.88 68.00 68.30
RegMean [19] 70.01 58.01 67.52 65.18
EMR-Merging [17] 74.20 67.11 70.98 70.76
Twin-Merging [24] 78.29 66.70 71.68 72.22
FREE-Merging(ours) 80.16 68.68 72.78 73.87
Table 3: Language model merging results: RoBERTa fine-tuned on 8 discriminative tasks, T0-3B on 11 discriminative tasks with $IA^{3}$, and Qwen-14B on 3 generative tasks with LoRA.
Method AlpacaEval GSM8K MBPP Avg.
LM Fine-tuned [41] 88.9 45.9 31.4 55.4
Math Fine-tuned [25] 22.3 63.5 23.0 36.5
Code Fine-tuned [26] 16.2 18.8 27.0 20.7
Task Arithmetic [18] 72.1 48.3 0 40.1
TIES-Merging [42] 77.5 66.9 27.2 57.2
DARE [47] 77.5 62.7 29.4 56.5
DELLA-Merging [7] 80.4 61.8 31.4 57.9
PCB-Merging [11] 81.2 60.6 30.7 57.5
FR-Merging(ours) 87.5 64.3 32.8 61.5
Twin-Merging [24] 83.5 61.4 30.4 58.4
EMR-Merging [17] 82.6 62.3 31.2 58.7
FREE-Merging(ours) 88.1 65.7 31.8 61.8
Table 4: Comparison of performance of different model merging methods after full fine-tuning LLaMa2-13B on three tasks: Instruction Following (LM), Math, and Code Generation.
Merging Large Language models. We investigate the merging of large models. We use LLaMa2-13B [35] as the pre-trained model, utilizing WizardLM [41] for instruction following, WizardMath [25] for math, and WizardCoder-Python [26] for coding. Experiments are conducted on the AlpacaEval (instruction following task), GSM8K (math task), and MBPP (code generation task) [47].
The results are shown in Table 4. It can be observed that our proposed FR-Merging even surpasses existing methods that rely on additional information, showing great advantages. This indicates that our findings in FR-Merging regarding low-frequency signals and task generalization also hold when extended to large models, which is highly significant for the practical deployment of large models. Additionally, our proposed FREE-Merging also shows considerable advantages. This suggests that FREE-Merging can also be effectively applied to large models, promoting rapid deployment of large models across cross-domain tasks.
Methods Task COCO-Retrieval COCO-Captioning ImageNet-1k Classification NLVR2 VQAv2
Metric Accuracy($\uparrow$) BLEU4($\uparrow$) CIDEr($\uparrow$) METEOR($\uparrow$) ROUGE-L($\uparrow$) Accuracy($\uparrow$) Accuracy($\uparrow$) Accuracy($\uparrow$)
Individual 0.8456 0.394 1.337 0.311 0.601 0.8537 0.7765 0.8439
Weight Averaging 0.1893 0.031 0.001 0.115 0.159 0.6771 0.2800 0.6285
Task Arithmetic [18] 0.3177 0.033 0.000 0.118 0.176 0.6732 0.3809 0.6933
Ties-Merging [42] 0.3929 0.029 0.001 0.108 0.167 0.6978 0.3206 0.6717
Breadcrumbs [6] 0.1893 0.011 0.012 0.065 0.114 0.6323 0.6073 0.6823
FR-Merging(ours) 0.7010 0.016 0.019 0.096 0.141 0.6992 0.6407 0.6965
EMR-Merging [17] 0.7946 0.289 1.060 0.272 0.534 0.7742 0.7475 0.7211
Twin-Merging [24] 0.7937 0.325 1.078 0.278 0.558 0.7730 0.7762 0.7232
FREE-Merging(ours) 0.8093 0.347 1.195 0.289 0.572 0.7850 0.8168 0.7297
Table 5: Results of merging multi-modal BEiT3 models on five vision-language tasks.
5.3 Merging Multi-Modal Models
We merge various fine-tuned models of the BEiT3-base [37] across five different tasks: ImageNet-1k (Image Classification), VQAv2 (Visual Question Answering), NLVR2 (Visual Reasoning), COCO Captioning (Image Captioning), and COCO Retrieval (Image-Text Retrieval). COCO Captioning is evaluated using BLEU4, CIDEr, METEOR, and ROUGE-L, while the other tasks are measured by accuracy [17]. Details are available in Appendix 9.1.
The results in Table 5 demonstrate that our method also shows significant improvements in multi-modal models. Unlike the merging of vision or language models, which often involve similar tasks, such as classification or generation, the task similarity here is relatively low. Therefore, the superiority in this configuration fully indicates the generalization ability of our FR-Merging and FREE-Merging, making them suitable for a wide range of applications.
Figure 3: Visualization of the loss landscape on EuroSAT before and after high-pass filtering (top vs bottom) in the model merging of the fine-tuned ViT-B/32 on EuroSAT and DTD tasks [18].
5.4 Fourier Transform Analysis
In this section, we explain why Fourier high-pass filtering improves merging performance from an optimization perspective. Previous work [21] shows that landscape geometry impacts model generalization. Fig 3 demonstrates that high-pass filtering widens the valley in the loss landscape (bottom plot). Since fine-tuned models from the same pre-trained model are closer in the landscape [30], a wider valley increases the likelihood of merged results falling within the loss basin as shown in the bottom plot of Fig 3, leading to better performance. Without filtering, the landscape is flat at a high level, which has little effect on model merging and only highlights the advantage of pre-training [15].
Ablation Part Avg. Acc(%)
Backbone Weight Averaging 65.80
+ FR-Merging 76.96 ($\uparrow$11.16)
+ $\lambda$ as Equa. 3 77.64 ($\uparrow$11.84)
Expert + Top-K Expert 79.38 ($\uparrow$13.58)
+ Our Rescaling 89.68 ($\uparrow$23.88)
Table 6: Ablation study of FREE-Merging. Gradually add each module proposed in this paper from top to bottom.
5.5 Ablation Study
In this section, we conduct ablation experiments on the components of our method. All experiments are performed on ViT-B/32 with eight classification tasks as target tasks.
Ablation on FR-Merging. In this part, we analyze the impact of replacing FR-Merging on the backbone performance and overall performance. Table 6 shows that adding FR-Merging leads to an 11.16% performance increase. Next, we replace high-pass filtering with other cost-free methods such as Ties [42], Top-K, Task Arithmetic [18], outlier dropout [6], EMR-Merging [17], and PCB-Merging [11], while also comparing low-pass filtering, and median filtering, with detailed results in Fig 4(a). To ensure a fair comparison, all experts use an additional 1% of parameters.
(a) Ablation on FR-Merging.
(b) Ablation on Task Expert.
(c) Router Analysis.
Figure 4: a) The ablation experiments on FR-Merging show that our proposed method achieves optimal performance. b) Performance of experts at different sparsity rates. c) In experiments with random routers, our approach demonstrates optimal performance.
Our proposed FR-Merging outperforms other cost-free methods on both the backbone and in combination with task experts, demonstrating its effectiveness in minimizing task conflicts in the merged network. Examining the different filtering methods, we find that low-pass filtering leads to great performance degradation, confirming that the low-frequency information contains task-specific signals and greatly affects generalization. Median-pass filtering also performs worse than FR-Merging, and this is because some high-frequency signals preserve abrupt changes in the signal and their removal harms the performance of the model.
Ablation on Merging Coefficients. This part presents an ablation study on our merging coefficients as in Eq. 3. Table 6 shows the results. It can be found that incorporating merging coefficients instead of simply averaging leads to a slight improvement. This is because the merging coefficients fully consider the variations of the outputs for each task before and after merging, ensuring that the output after merging is similar to the one before merging. This is especially suitable for cases where the magnitudes of the fine-tuning parameters of different tasks are inconsistent.
Ablation on Task Expert. We conduct an ablation study on the expert extraction method, First, Table 6 shows that incorporating our proposed experts and rescaling method leads to great improvements, confirming the effectiveness.
To validate the superiority of our method, we evaluate the performance of only utilizing experts while retaining different proportions of parameters. We compare the original Top-K selection, Bernoulli selection [47], SVD extraction [24], and our proposed Top-K extraction with rescaling. As shown in Fig 4(b), our method shows substantial advantages, particularly under high sparsity conditions, supporting the effectiveness of our lightweight expert extraction.
5.6 Router Analysis
In this part, we analyze the proposed router method. We aim to achieve strong performance even if the router performs sub-optimally. Router inaccuracies are likely to occur due to insufficient training resources or issues during the training on edge deployment. Therefore, the advantages of our methods in this scenario are of great practical significance.
We use a random router to simulate an inaccurate router and compare our method with several existing router-based methods. Our proposed method outperforms Twin-Merging by 13.5% and outperforms EMR-Merging by 8.5% on average, as shown in Fig 4(c), demonstrating that our method constructs a reasonable backbone and generalizable experts.
5.7 Speed and Storage Space Analysis
Method Training Tokens Training Cost Storage Param. Inference Cost (/1000 items) Acc.
Individual 0 0 2.4B 31.0s 94.2
MTL 304M 12h 304M 31.0s 93.5
FR-Merging 0 0 304M 31.0s 88.2
Twin-Merging 0.9M 147s 352M 32.2s 92.7
FREE-Merging 0.9M 147s 328M 32.2s 93.7
Table 7: Trade-off of performance, training, inference, and storage costs using ViT-L/14 across 8 image classification tasks.
This part analyzes the training costs, inference speed, and storage requirements; details are shown in Table 7. Our FR-Merging achieves strong performance with a single model for storage, significantly reducing storage demands without extra training or inference costs. FREE-Merging, with a small increase in training and storage overhead, further enhances performance. These two methods greatly facilitate the practical deployment of deep learning at minimal cost.
6 Conclusion
In this paper, we propose FR-Merging and FREE-Merging to efficiently merge models, leveraging existing models to create a multi-task model. FR-Merging builds a backbone without training by high-pass filtering to remove information that hinders generalization, while FREE-Merging employs lightweight experts to offset information loss during merging. Our method balances training, storage, inference costs, and performance, showing high practical value.
Supplementary Material
7 Notations
We first list the notations for key concepts in our paper.
Notations Descriptions
$f_{\theta}$ A neural network with parameters $\theta$.
$\theta_{pre}$ Pre-trained model parameters.
$\theta_{k}$ Fine-tuning parameters for task $k$.
$\theta_{m}$ Model merging backbone network.
$v_{k}$ Task vector for task $k$, $v_{k}=\theta_{k}-\theta_{pre}$.
$M(v_{k},d)$ The top 100dparameters with the most significant changes corresponding to $v_{k}$.
$\mu_{k}$ The rescaling coefficient of the expert of task $k$
$e_{k}$ Lightweight task expert for task $k$, $e_{k}=\mu_{k}M(v_{k},d)$.
$\theta_{*}$ The backbone with the corresponding expert.
$G(\cdot)$ Transformation operations on parameters.
$\lambda_{k}$ Merging coefficient of the model parameters for task $k$.
$x\in R^{m}$ The input of the neural network.
$Y\in R^{n}$ The output of the neural network.
Table 8: Notations.
8 Proof
In this section, we provide a proof for Theorem 4.1. We begin by introducing the notation used in this proof.
8.1 Notations and preliminaries
To illustrate the proof, let’s consider two fine-tuned models with parameters $\theta_{A}$ and $\theta_{B}$, For clarity, we focus on a single-layer MLP (without activation functions), though we will explain at the end how our proof generalizes to all cases. The merged result is given by: $\theta_{m}=\lambda_{1}G(\theta_{A})+\lambda_{2}G(\theta_{B})$, where $G(\cdot)$ represents a transformation operation applied to the parameters. Assume that $\theta_{A}$ corresponds to task input $X_{A}$, and $\theta_{B}$ corresponds to task input $X_{B}$.
The goal is to prove that, without additional storage or extra training for $\theta_{A}$ and $\theta_{B}$, simply storing $\theta_{m}$ cannot perfectly retain the capabilities of both $\theta_{A}$ and $\theta_{B}$.
8.2 Details of proof
Next, we will analyze the cases separately.
Without transformations. Assume that we do not apply transformations to the output of the merged network. the output of the model $f_{\theta}$ can be expressed as:
$output_{f}=\theta x$ (5)
Suppose $\theta_{m}$ can perfectly retain the capabilities of both $\theta_{A}$ and $\theta_{B}$ without introducing new knowledge. That is, for $\forall x_{A}\in X_{A},x_{B}\in X_{B}$, we have:
$\theta_{m}x_{A}=\theta_{A}x_{A},\ \theta_{m}x_{B}=\theta_{B}x_{B}.$ (6)
Then we have:
$[\lambda_{1}G(\theta_{A})+\lambda_{2}G(\theta_{B})]x_{A}=\theta_{A}x_{A},$ (7)
$[\lambda_{1}G(\theta_{A})+\lambda_{2}G(\theta_{B})]x_{B}=\theta_{B}x_{B}.$ (8)
Assume that $\exists\ x_{m}=\mu_{1}x_{A}+\mu_{2}x_{B}\in X_{A}$, where $\mu_{1},\mu_{2}\in\mathbb{N}$. Substituting, we can obtain:
$\theta_{m}x_{m}=\theta_{A}x_{m}$ (9)
$\Rightarrow\theta_{m}(x_{A}-x_{m})=\theta_{A}(x_{A}-x_{m})$ (10)
$\Rightarrow\theta_{m}[(1-\mu_{1})x_{A}-\mu_{2}x_{B}]$ (11)
$\quad=\theta_{A}[(1-\mu_{1})x_{A}-\mu_{2}x_{B}]$ (12)
$\Rightarrow\mu_{2}\theta_{m}x_{B}=\mu_{2}\theta_{A}x_{B}$ (13)
$\Rightarrow\mu_{2}\theta_{B}x_{B}=\mu_{2}\theta_{A}x_{B}$ (14)
Due to:
$\mu_{2}\theta_{m}x_{B}=\mu_{2}\theta_{B}x_{B}$ (15)
Thus, if the Eq. 14 and Eq. 15 holds for $\forall x_{A}\in X_{A},x_{B}\in X_{B}$, then $\mu_{2}=0$. Therefore, we have:
$\forall\mu_{2}\neq 0,\nexists\ x_{m}=\mu_{1}x_{A}+\mu_{2}x_{B}\in X_{A}$ (16)
Eq. 16 indicates that $X_{A}$ is not continuous. However, we usually consider the input space of the model to be continuous. Even if it is discontinuous, we expect the model to be robust to slight perturbations in the input.
Thus, the assumption that $\theta_{m}$ can perfectly retain $\theta_{A}$ and $\theta_{B}$ without extra data or training does not hold in this case.
With transformations. Assuming we apply certain transformations to the output of the merged network, the output of the model $f_{\theta}$ can be expressed as:
$output_{f}=h(\theta x),$ (17)
where $h(\cdot)$represents a transformation applied to the output.
By the universal approximation theorem, $h(\cdot)$ can be approximated by a two-layer MLP with sigmoid activation. Then, assuming $\theta_{m}$ can perfectly retain the capabilities of both $\theta_{A}$ and $\theta_{B}$. Then for $\forall x_{A}\in X_{A},x_{B}\in X_{B}$, we have:
$W_{2}[sigmoid(W_{1}\theta_{m}x_{A})]=\theta_{A}x_{A}$ (18)
$W_{2}[sigmoid(W_{1}\theta_{m}x_{B})]=\theta_{B}x_{B}.$ (19)
Substituting, we obtain:
$\frac{W_{2}}{1+e^{-W_{1}\theta_{m}x_{A}}}=\theta_{A}x_{A},$ (20)
$\frac{W_{2}}{1+e^{-W_{1}\theta_{m}x_{B}}}=\theta_{B}x_{B}.$ (21)
Then we have:
$W_{2}$ $=\theta_{A}x_{A}(1+e^{-W_{1}\theta_{m}x_{A}})$ (22)
$=\theta_{B}x_{B}(1+e^{-W_{1}\theta_{m}x_{B}}).$ (23)
Thus, we can obtain the following equation:
$\theta_{A}x_{A}(e^{W_{1}\theta_{m}x_{A}}+1)e^{W_{1}\theta_{m}x_{B}}$ (24)
$=\theta_{B}x_{B}(e^{W_{1}\theta_{m}x_{B}}+1)e^{W1\theta_{m}x_{A}}.$ (25)
Therefore,
$(\theta_{A}x_{A}-\theta_{A}x_{B})(e^{W_{1}\theta_{m}x_{B}}e^{W_{1}\theta_{m}x_{A}}+e^{W_{1}\theta_{m}x_{B}})$ (26)
$=\theta_{B}x_{B}(e^{W_{1}\theta_{m}x_{B}}+e^{W_{1}\theta_{m}x_{A}})$ (27)
Therefore, based on Equa. 22 and Equa. 26, since they hold for $\forall x_{A}\in X_{A},x_{B}\in X_{B}$, at least one of $W_{1}$ or $W_{2}$ can be expressed as $\Phi(\theta_{A},\theta_{B})$, where $\Phi$ represents a function.
In other words, only by additionally storing part of the information from $\theta_{A}$ or $\theta_{B}$ and processing $\theta_{m}$ at the output can $\theta_{m}$ perfectly retain the capabilities of both $\theta_{A}$ and $\theta_{B}$. Therefore, the assumption that $\theta_{m}$ can perfectly retain both $\theta_{A}$ and $\theta_{B}$ does not hold in this case.
Next, we briefly outline the rationale for extending this result to all layers of a neural network. Our assumptions involve only a single layer of the neural network, which already cannot perfectly retain capabilities without introducing new data. Therefore, it is even less feasible for the entire network to achieve this. Thus, this extension is reasonable.
In summary, the proposition is proven.
9 Reproducibility
9.1 Datasets
Merging 8 ViTs. We use ViT-B/32 and ViT-L/14 as pre-trained models, and fine-tune them on 8 image classification datasets (SUN397, Cars, RESISC45, EuroSAT, SHVN, GTSRB, MNIST, and DTD), then merge the models and test their performance. Configuration details follow [17].
Merging 30 ViTs. We use ViT-B/16 as the pre-trained model and test its merging performance on 30 image datasets. The datasets include MNIST, CIFAR-10, Vegetables, Food-101, Kvasir-v2, Cars, Intel Images, EuroSAT, Weather, Cats and Dogs, MangoLeafBD, Beans, CIFAR-100, GTSRB, SHVN, Dogs, Fashion MNIST, Oxford-IIIT-Pet, Landscape Recognition, Flowers Recognition, STL-10, CUB-200-2011, EMNIST, DTD, RESISC45, SUN397, KenyanFood13, Animal-10N, Garbage Classification, and Fruits-360. These datasets cover a wide range of major image categories. Configuration details follow [17].
Merging Medium-sized Language Models. We use RoBERTa as the pre-trained model, fine-tune it on 8 classification task datasets from GLUE benchmark for model merging, including CoLA, SST-2, MRPC, STS-B, QQP, MNLI, QNLI, and RTE. CoLA is evaluated with the Matthews correlation coefficient, STS-B with the average of the Pearson and Spearman correlation coefficients, and the others by accuracy. Details follow [17].
Merging PEFT Models. In this section, we conduct two experiments. The first uses T0-3B as the pre-trained model and $IA^{3}$ as the PEFT method, fine-tuned and merged on RTE, CB, Winogrande, WiC, WSC, COPA, H-SWAG, Story Cloze, and ANLI (R1 to R3). Details can be found in [17]. The second part uses Qwen-14B as the pre-trained model and LoRA as the PEFT method. We fine-tune and merging on three generative tasks: MMLU, TruthfulQA, and BBQ. Configuration details can be found in [24].
Merging Large Language Models. We use LLaMa2-13B as the pre-trained model and apply fine-tuned results from WizardLM, WizardMath, and WizardCoder-Python for instruction following, math, and code tasks, respectively. We perform merging and test the model on three datasets: AlpacaEval (instruction following task), GSM8K (math task), and MBPP (code generation task). Details follow [47].
Merging Multi-Modal Models. We use BEiT3 as the pre-trained model, fine-tune it on five multi-modal datasets, and test model merging performance. The datasets are ImageNet-1k (Image Classification), VQAv2 (Visual Question Answering), NLVR2 (Visual Reasoning), COCO Captioning (Image Captioning), and COCO Retrieval (Image-Text Retrieval). COCO Captioning is evaluated using BLEU4, CIDEr, METEOR, and ROUGE-L, while the other tasks are measured by accuracy. Details follow [17].
9.2 Baselines
Individual. Solve each task with its fine-tuned model, but this requires storing separate models for each task, leading to significant storage overhead.
Traditional MTL. Train a model on data from all tasks, which incurs high training costs and data privacy issues.
Weight Averaging. Average the parameters for merging. The method is simple, but it faces great performance loss.
Fisher Merging [28]. Use Fisher information matrices to assess parameter importance and determine merging coefficients. However, this requires calculating the matrix for each model, demanding high computational resources, making it unsuitable for edge deployment.
Task Arithmetic [18]. Define task vectors as the merging target. For task $k$, the task vector is defined as $v_{k}=\theta_{k}-\theta_{pre}$, where $\theta_{pre}$ is the pre-trained model parameters, and $\theta_{k}$ is the fine-tuned parameters for task $k$. The merging process can be represented as $\theta_{m}=\theta_{pre}+\lambda\sum_{i=1}^{K}v_{i}$, where $\lambda$ is the merging coefficient. This method suffers significant performance degradation due to unaddressed task conflicts.
Ties-Merging [42]. Attempts to resolve parameter conflicts during model merging by eliminating redundancy and sign conflicts. However, resolving parameter conflicts is insufficient to address task conflicts, resulting in performance loss.
Breadcrumbs [6]. Discards parameters with the largest and smallest absolute values as redundant, negatively impacting model merging. This approach is simple, but it shows a significant performance decline on certain tasks.
PCB-Merging [11]. Uses internal balancing to measure parameter importance within tasks and mutual balancing to assess parameter similarity across tasks, discarding redundant parameters and adjusting merging coefficients. This approach requires considerable computational resources, whereas our FR-Merging approach requires less computational power while yielding better results.
RegMean [19]. A weighted merging model based on a closed-form solution for the merging problem, aiming to maximize the similarity between the merged model and each pre-merged model on the same input. Requires inner product data from training. Since this method relies on data information, which is typically not provided by the model provider, its applicability is limited to specific scenarios.
AdaMerging [46]. Uses an unsupervised approach to learn the merging coefficient for each task vector or layer. AdaMerging++ additionally applies Ties-Merging before calculating the merging coefficient. This method is limited to classification tasks and requires certain training resources, making it unsuitable for edge deployment.
DARE [47]. DARE randomly discards a large portion of task vector parameters before merging, potentially reducing parameter interference among models. This method is simple, but due to the lack of further optimization in the merging, it suffers from significant performance degradation.
EMR-Merging [17]. Proposes retaining a mask matrix and rescale parameters for each model. During inference, the mask matrix and rescaling parameters are selected to recover performance. This method cannot further improve performance because it does not optimize the merged backbone. Additionally, since a mask must be stored for each task, it requires considerable storage space. In contrast, our method optimizes these aspects simultaneously.
Twin-Merging [24]. Proposes saving an expert module extracted by SVD from the original parameters for each task, which is dynamically added during inference. This method does not optimize the merged backbone, resulting in suboptimal performance. It requires storing a large number of parameters when constructing task experts. In contrast, our proposed method optimizes both aspects, making it suitable for resource-constrained edge deployment scenarios.
9.3 Details
In this section, we discuss the details of the experiments. All experiments, except for the speed tests, are conducted on eight NVIDIA RTX-3090 GPUs. The inference speed tests are performed using a single RTX-3090 GPU.
10 More Discussions and Future Directions
10.1 Role of Task Experts
In this section, we demonstrate that introducing additional information can potentially mitigate task conflicts.
We use ViT-B/32 as the pre-trained model and fine-tune it on 8 image classification tasks, following the setup in [18]. To prevent the impact of parameter erasure, we select only 5% of non-overlapping parameters for each task. As a result, after this selection, only task conflicts affect the model merging process. Merging resulted in a significant performance drop, indicating the presence of task conflicts. However, when we double task-specific vectors for each task during inference, performance improves substantially. This suggests that introducing additional information for each task has the potential to resolve task conflicts.
Avg acc(%) Norm acc(%)
Fine-tuned 90.5 100.0
No Overlap 5% params. 89.91 99.34
No Overlap Merging 67.31 74.37
+ Task Knowledge 84.55 93.42
Table 9: Using ViT-B/32 as the pre-trained model, we fine-tune on eight classification tasks, selecting 5% non-overlapping parameters per task. Performance improves significantly by adding task-specific parameters during inference.
10.2 Task Experts Construction Analysis
Figure 5: The analysis evaluates the reasonableness of approximating task-specific information using the most changed parameters during fine-tuning, by calculating the L1 distance of each layer’s parameters in CLIP-B/32 fine-tuned on 8 visual tasks with different approximation methods. Here, Ori refers to the original parameters, FP to low-pass filtering, Top to maximum value approximation, and Rand to random sampling approximation.
This section discusses the rationale behind using the parameters with the largest changes during the fine-tuning to approximate task-specific information introduced by fine-tuning, thereby constructing task experts.By comparing L1 distances of model parameters in the frequency domain, with smaller distances indicating better approximation, we assess the effectiveness of different methods.
Based on the analysis in Sec 4.3, low-frequency information contains more task-specific information but is discarded during model merging in our proposed FR-Merging. Therefore, task experts should compensate for this missing information. First, we briefly explain why approximation methods are preferred over directly saving low-frequency information. If low-frequency data is directly saved, ensuring lightweight storage requires saving frequency-domain data instead of inverse-transformed results. This would necessitate performing inverse transformations during expert invocation, significantly increasing the inference time. Considering the constraints of time and storage in edge-side deployment, a more efficient approximation method is needed.
Next, we compare the extent to which different approximation methods approximate low-frequency information to assess their suitability for constructing task experts. In Fig 5, we present the effects of approximating using the top 5% largest fine-tuned parameters and random sampling. It can be observed that using the parameters with the largest changes during the fine-tuning closely approximates the frequency-domain characteristics of low-frequency information, strongly supporting the rationale for preserving task-specific information by saving the parameters with the largest changes during the fine-tuning. Additionally, the high similarity between low-frequency information and the original information confirms our hypothesis: low-frequency information contains more task-specific information introduced during fine-tuning.
In summary, using the parameters with the largest changes during fine-tuning as task experts to offset task-specific information lost during merging is reasonable. Furthermore, as validated in Sec 5.5, employing our proposed rescaling method yields even better results. Thus, the expert construction method proposed in this paper is justified.
10.3 Filtering Frequency Analysis
Filtering Percentage SUN397 Cars RESISC45 EuroSAT SVHN GTSRB MNIST DTD Avg.
Individual 75.3 77.7 96.1 99.7 97.5 98.7 99.7 79.4 90.5
0% 65.3 63.4 71.4 71.7 64.2 52.8 87.5 50.1 65.8
10% 66.1 65.3 77.2 90.0 84.3 81.0 98.2 58.4 77.6
20% 65.6 64.5 76.9 90.3 84.6 82.6 98.5 59.1 77.7
30% 66.2 64.5 77.2 90.1 85.4 82.3 98.5 60.0 78.1
40% 65.2 64.4 76.1 90.4 84.5 82.2 98.5 59.0 77.5
50% 64.9 64.4 76.3 88.9 82.3 81.2 98.4 58.9 76.9
60% 65.4 65.0 76.9 86.7 77.3 76.6 96.9 57.7 75.3
Table 10: The performance of Fourier high-pass filtering at different filtering percentages.
In this section, we analyze the cutoff frequency of the Fourier high-pass filter. We perform merging using the results after fine-tuning the ViT-B/32 on eight visual classification tasks. We explore the impact of filtering different proportions of low-frequency signals on model merging.
The results are shown in the Table 10. It can be observed that filtering a certain proportion of low-frequency information significantly improves the performance of the model merging. For example, filtering 10% of the low-frequency information, despite its small proportion, leads to an impressive 12% improvement in average performance. This is similar to the analysis of low-frequency information in image processing, where low-frequency components contain most of the features in the overall information. In the context of model merging, low-frequency information represents the specialized knowledge brought by fine-tuning. While this specialized information can improve model performance on specific tasks, it severely affects the generalization ability of the model, leading to task conflicts during merging. Therefore, filtering a small proportion of low-frequency information helps minimize model specialization without significantly impacting performance, thus avoiding performance loss caused by task conflicts.
From the experimental results in Table 10, it is evident that our method is robust to the filtering ratio. Even after filtering 40% of the information, the method still demonstrates strong performance improvements, with a noticeable performance drop only after filtering 60% of the information. This strongly indicates that the method proposed in this paper is highly robust to the cutoff frequency as a hyper-parameter. The robustness to hyper-parameters makes our method simple to deploy, providing a significant advantage.
10.4 Future Works
In this section, we outline directions for our future work. First, we aim to extend our proposed model merging approach to fine-tuned models with different architectures, enabling broader applicability. Additionally, we plan to further investigate the differences between PEFT and full fine-tuning and design customized model merging methods tailored to PEFT, enabling effective utilization and deployment of existing PEFT results.
11 Additional Experimental Results
11.1 Merging Vision Models
In this section, we present the results of merging using ViT-L/14 as the pre-trained model across the 30 vision datasets mentioned in Sec. 5.1, as shown in Table 11.
From the results, we can see that, consistent with the conclusions in the main text, our FR-Merging outperforms all cost-free methods, efficiently constructing a high-performance merged backbone. With the dynamic addition of experts, FREE-Merging achieves satisfactory performance at a minimal additional cost.
Task-specific Acc MNIST Cifar-10 Vegetables Food-101 Kvasir-v2 Intel-Images Cars EuroSAT Weather Cats and Dogs
Individual 99.22 97.88 100.00 87.93 94.31 94.63 85.96 99.04 98.22 99.05
Weight Averaging 27.63 42.91 83.20 68.02 25.27 82.40 7.74 24.37 61.06 91.28
Task Arithmetic 30.81 59.86 91.97 73.06 31.05 89.03 9.34 31.25 74.56 93.61
Ties-Merging 23.21 42.82 92.31 73.22 21.09 89.39 5.30 10.98 72.86 91.88
Breadcrumbs 33.95 54.29 91.83 66.24 32.12 14.30 0.19 32.33 69.04 90.92
FR-Merging(ours) 42.78 63.35 93.16 76.25 38.13 90.73 12.00 36.37 75.60 96.69
RegMean 90.71 89.65 99.10 76.14 71.00 93.60 16.28 74.13 86.62 98.54
AdaMerging 81.22 87.54 97.97 75.23 22.76 91.02 0.42 44.60 89.13 96.91
Twin-Merging 92.10 89.34 98.54 81.43 78.45 94.14 58.12 69.43 92.23 98.85
EMR-Merging 93.40 90.23 98.34 83.54 76.54 95.12 54.34 76.64 94.54 98.65
FREE-Merging(ours) 97.30 95.60 99.00 84.65 85.25 94.45 58.52 79.12 96.38 98.62
Dogs Fashion Pet LandScape Flowers STL-10 CUB-200-2011 EMNIST DTD RESISC45
Individual 85.16 93.26 92.23 86.83 98.19 99.07 84.79 94.67 71.76 98.90
Weight Averaging 47.80 20.46 31.26 73.14 68.97 37.74 37.66 7.73 14.63 13.56
Task Arithmetic 47.65 37.11 33.24 79.59 80.68 39.66 41.86 11.05 14.73 15.50
Ties-Merging 26.03 27.05 12.84 78.27 34.33 6.17 31.28 5.61 3.71 6.79
Breadcrumbs 40.91 35.29 39.16 83.80 81.94 54.73 27.58 3.25 17.81 15.20
FR-Merging(ours) 49.86 42.73 41.26 85.80 84.49 59.53 41.79 23.80 22.23 19.49
RegMean 42.89 83.42 34.62 83.64 95.26 78.94 49.78 48.67 30.53 34.66
AdaMerging 53.09 76.76 48.34 81.98 95.69 68.91 48.19 18.02 16.68 24.83
Twin-Merging 72.65 80.64 68.78 85.49 96.75 74.72 60.47 57.47 41.25 50.57
EMR-Merging 72.34 82.35 67.35 85.73 96.46 76.65 63.11 60.03 40.48 49.57
FREE-Merging(ours) 75.86 86.43 72.57 86.42 97.54 79.01 65.65 63.64 45.71 54.25
MangoLeafBD Beans Cifar-100 GTSRB SVHN SUN397 KenyanFood13 Animal-10N Garbage Fruits-360
Individual 100.00 97.73 89.85 95.74 96.22 78.98 85.53 92.52 93.36 99.63
Weight Averaging 68.58 70.98 77.98 15.00 10.88 57.42 33.55 46.00 22.89 5.38
Task Arithmetic 87.02 84.62 80.20 37.01 17.41 55.88 36.32 51.14 25.23 6.15
Ties-Merging 76.58 67.22 78.61 40.74 10.54 52.69 19.90 19.13 3.91 1.50
Breadcrumbs 79.25 85.15 75.65 38.51 22.89 47.96 26.07 17.303 28.06 1.40
FR-Merging(ours) 86.50 87.38 82.03 44.19 30.08 56.98 38.90 49.10 32.01 12.00
RegMean 98.10 92.58 82.59 56.96 66.13 58.58 57.11 68.74 65.31 19.79
AdaMerging 99.13 93.38 84.19 59.90 25.70 64.09 48.66 66.55 38.54 7.94
Twin-Merging 100.0 94.24 85.57 65.87 75.49 71.25 63.81 71.67 70.82 25.56
EMR-Merging 100.0 95.45 86.43 65.94 76.46 72.45 65.43 73.84 72.24 28.24
FREE-Merging(ours) 100.0 96.40 88.65 68.63 82.23 76.78 69.45 78.40 75.84 37.97
Average Acc Individual Weight Averaging Task Arithmetic Ties-Merging Breadcrumbs FR-Merging(ours)
Acc 93.02 42.52 48.89 37.53 43.57 53.91
Average Acc Individual RegMean AdaMerging Twin-Merging EMR-Merging FREE-Merging(ours)
Acc 93.02 68.14 60.25 75.52 76.39 79.67
Table 11: Task-specific and average performance when merging ViT-B/16 models on 30 tasks.
11.2 Merging Language Models
Method CoLA SST2 MRPC STSB QQP MNLI QNLI RTE Avg
Individual 60.18 94.04 89.22 90.63 91.41 87.20 92.71 79.06 85.55
Weight Averaging 13.96 64.11 69.36 31.84 75.36 42.19 58.70 55.23 51.34
Task Arithmetic 18.78 85.89 79.90 74.03 83.78 59.08 69.67 62.09 66.65
Ties-Merging 20.48 84.40 81.13 58.19 85.70 64.65 74.81 42.96 64.04
Breadcrumbs 25.23 85.21 78.19 68.45 81.07 64.40 79.83 63.18 68.19
PCB-Merging 23.46 84.59 79.36 68.89 81.98 63.57 76.56 64.47 67.86
FR-Merging(ours) 15.25 86.81 82.39 80.81 86.21 65.69 78.26 64.73 70.02
RegMean 36.67 90.60 75.74 62.68 83.55 70.02 82.35 58.48 70.01
EMR-Merging 37.82 87.91 82.31 72.90 82.61 78.91 83.24 67.92 74.20
Twin-Merging 52.30 93.23 88.65 68.64 81.99 78.80 88.40 74.37 78.29
FREE-Merging(ours) 54.50 93.69 88.46 67.04 88.03 80.60 89.90 79.06 80.16
Table 12: Results of merging RoBERTa pre-trained models on eight datasets from GLUE benchmark.
Method RTE CB Winogrande WiC WSC COPA H-SWAG Story Cloze ANLI-R1 ANLI-R2 ANLI-R3 Avg. Acc
Individual 82.7 95.8 75.1 71.7 65.3 85.3 44.4 94.9 70.2 46.5 53 71.4
Traditional MTL 88.6 95.8 75.5 61.1 80.6 94.1 42.3 97.6 70.5 49.8 47.7 73.1
Fisher Merging 83.3 83.3 56.7 54.2 58.3 83.1 42.2 94.1 45.9 41.0 42.2 62.2
Task Arithmetic 74.1 83.3 62.8 49.1 49.3 87.5 41.5 95.3 60.8 49.4 50.0 63.9
Weight Averaging 81.2 58.3 53.8 55.2 53.5 80.9 40.1 92.5 43.3 39.2 40.2 58.0
Ties-Merging 78.0 83.3 67.9 57.6 59.7 81.7 42.8 90.3 66.9 51.3 51.1 66.4
Breadcrumbs 71.9 59.4 53.1 31.2 64.1 78.1 46.9 90.2 50.0 31.2 25.0 54.6
PCB-Merging 85.9 83.3 61.9 57.1 63.9 82.4 42.7 91.2 64.2 47.8 45.9 66.1
FR-Merging(ours) 81.3 83.8 68.3 57.9 65.2 88.4 48.6 93.5 56.2 46.2 46.3 66.9
RegMean 81.2 58.3 53.8 55.2 53.5 80.9 40.1 92.5 43.3 39.2 40.2 58.0
EMR-Merging 81.8 87.5 66.6 56.1 65.3 82.4 44.7 93.6 65.7 43.8 50.8 67.1
Twin-Merging 81.2 85.6 65.4 57.2 66.3 81.6 44.4 92.9 66.5 42.4 51.2 66.8
FREE-Merging(ours) 83.2 89.3 69.9 58.4 65.5 84.3 45.2 94.2 67.9 45.2 52.4 68.7
Table 13: Results of merging $(IA)^{3}$ models with T0-3B as pre-trained model on eleven NLP tasks.
Method MMLU TruthfulQA BBQ Avg.
Pre-trained 69.30 51.27 80.69 67.09
Fine-tuned 68.35 54.34 93.53 72.07
Weight Averaging 68.10 50.01 82.31 66.80
Task Arithmetic 67.62 53.38 78.24 66.41
Task Arithmetic (w/ DARE) 67.82 52.66 82.83 67.77
Ties-Merging 68.27 50.01 84.10 67.46
Ties-Merging (w/ DARE) 69.32 53.07 81.19 67.86
Breadcrumbs 68.24 52.88 79.20 66.77
PCB-Merging 67.23 52.48 80.37 66.69
FR-Merging(ours) 68.16 53.39 82.44 68.00
RegMean 67.89 52.45 82.24 67.52
EMR-Merging 67.82 55.01 90.13 70.98
Twin-Merging 68.32 55.76 90.98 71.68
FREE-Merging(ours) 68.83 57.39 92.14 72.78
Table 14: Results of merging LoRA models with QWEN-14B as pre-trained model on three generative tasks.
In this section, we present the complete experimental results of merging language models.
First, we show the merging results for RoBERTa as the pre-trained model, fine-tuned separately on the eight tasks of GLUE, as summarized in Table 12. Our method achieves optimal results on nearly all datasets, demonstrating its generalization capability on language models. Additionally, Table 13 presents the merging performance on 11 language classification tasks using T0-3B as the pre-trained model and $IA^{3}$ as the fine-tuning method. Our approach similarly shows an advantage here. Finally, Table 14 displays the merging performance for QWEN-14B as the pre-trained model and LoRA as the fine-tuning method across three generative tasks. Both our FR-Merging and FREE-Merging exhibit significant advantages over competing methods in each respective category.
Overall, our proposed methods demonstrate strong generalization capabilities on language models.